OpenStudy (dschneider2016):

Just check my answer?!!! Will medal!! Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

OpenStudy (dschneider2016):

trig identities for sum and difference: cos (a + b) = cos a cos b - sin a sin b cos (a - b) = cos a cos b + sin a sin b cos (x - y) - cos (x + y) = cos x cos y + sin x sin y - (cos x cos y - sin x sin y) = cos x cos y + sin x sin y - cos x cos y + sin x sin y = 2 sin x sin y

OpenStudy (dschneider2016):

@Kkutie7

zepdrix (zepdrix):

So this last step here, you did what? `cos x cos y + sin x sin y - cos x cos y + sin x sin y` You just combined like-terms?

zepdrix (zepdrix):

seems good :)

OpenStudy (dschneider2016):

So should I change anything? @zepdrix

zepdrix (zepdrix):

nah, seems perfect

OpenStudy (dschneider2016):

Thanks so much!!